Calculate the velocity of an electron placed in the 2nd orbit of Li2+ ion.Also
Calculate the number ofrevolutions per second that it makes around the nucleus.

Grade:11

1 Answers

vikas askiitian expert
13 years ago

V=Vo(Z/n)               ( Z is atomic number and n is principal qualtam number)

 Vo=2* 10^6m/s

  V=Vo3/2          (Z for li is 3)

V=9*10^6m/s................1

V=wR 

R( radius) for nth orbit is R=Ro(n^2/Z)

  Ro=52.9pm=5.29*10^-11m

R(for second orbit) =7.06*10^-11m

w = v/r

   =1.27*10^17 revolutions/sec........2